solve the system of equations 3y-2x=4 and 2y+4x=8
Answer by Evy(14)

Hello, the first thing you need to do is:
Multiply the first equation by 2
2(3y-2x=4) which gives 6y-4x=8


then solve:
6y-4x=8
2y+4x=8
4x cancels out


You now have:
6y=8
2y=8
-------- Add like terms
8y=16


Now divide by 8


8y = 16
--------
8


y=8



to check you answer just plug in the numbers.


If you have any questions you can e-mail me Southqueensb@yahoo.com
~EVY~